What voltage mist be applied to an xray tube for it to emit xrays with a minimum wavelength of 30picometres


1
Expert's answer
2021-03-17T16:55:40-0400
"qV=\dfrac{hc}{\lambda},""V=\dfrac{hc}{q\lambda},""V=\dfrac{6.62\cdot10^{-34}\ J\cdot s\cdot3\cdot10^8\ \dfrac{m}{s}}{1.6\cdot10^{-19}\ C\cdot30\cdot10^{-12}\ m}=4.14\cdot10^4\ V."